Sweetie145:

Solve the linear equation `2(3x+2)=2x-1+x`.  State the property that justifies your first step and why you chose it.

Extrinix:

So to start off, let's use the `distributive property` and simplify the equation \(2(3x+2)=2x-1+x\) \(6x+4=2x-1+x\) Now we can solve for \(x\) \(6x+4=2x-1+x\) Combine like terms \(6x+4=3x-1\) Subtract 3x from both sides \(3x+4=-1\) Subtract 4 from both sides \(3x=-5\) Divide by 3 on both sides, and you have \(x\) \(x=-\dfrac{5}{3}\)

Sweetie145:

@extrinix wrote:
So to start off, let's use the `distributive property` and simplify the equation \(2(3x+2)=2x-1+x\) \(6x+4=2x-1+x\) Now we can solve for \(x\) \(6x+4=2x-1+x\) Combine like terms \(6x+4=3x-1\) Subtract 3x from both sides \(3x+4=-1\) Subtract 4 from both sides \(3x=-5\) Divide by 3 on both sides, and you have \(x\) \(x=-\dfrac{5}{3}\)
omg thank u so much 4 helping me 🥰💙✨